A force applict tody Z by changed on a mass 16 relacity from 10 m/s to 54 myys. During 18 body covered a distan 8 its this interal the 10 m. Find the force applied.


1
Expert's answer
2021-07-05T08:42:59-0400

A force applied to body Z of a mass 16 kg changed its velocity from 10 m/s to 54 m/s. During this process, the body covered a distance of 10 m. Find the force applied.


The force is


"F=ma,\\\space\\
a=\frac{v_f^2-v_i^2}{2d},\\\space\\
F=m·\frac{v_f^2-v_i^2}{2d}=2253\text{ N}."